OpenStudy (anonymous):

Find the lateral area and total area for the prism. Enter the answers below.

OpenStudy (karatechopper):

where is the prism

OpenStudy (anonymous):

Well you can see it if you look at the doc... let me know if you need a different kind of file to open it.

OpenStudy (stormfire1):

From the picture, you can see that it's a right prism and as you probably know, the lateral area is that of all the sides excluding the 2 bases. So the lateral area = (6*3)+(6*3)+6*3) = 54 sqin To find the total area you need to also find the area of the 2 bases. Since the area of a triangle is 1/2 base * height the area of one of the faces will be 1.5 * 3 = 4.5 sqin. Since there are 2 bases double that to get 9 sqin. So the final answer will be 54 + 9 = 63 sq in.